OpenStudy (anonymous):

If a flea can jump straight up to a height of 21.1 , what is its initial speed as it leaves the ground, neglecting air resistance?

OpenStudy (anonymous):

what are the units of 21.1 ?

OpenStudy (anonymous):

basically the concept underlying here is the conservation of energy. As only conservative force (gravitational force) acts, total mechanical energy is conserved. \[mgh = 1/2 mv ^{2}\] \[ v=\sqrt{2gh}\] where 'h' is the height to which the flea jumped.